A thin spherical shell of radius 6 . 68 m has a total charge of 9 . 17 C distributed uniformly over its surface. a) Find the electric field 11 . 3 m from the cen- ter of the shell. The Coulomb constant is 8 . 98755 � 10 9 N � m 2 / C 2 . Answer in units of N / C
b)Find the electric field 1 . 53 m from the center of the shell. Answer in units of N / C
Answer #1

1 ) k*q/r^2= 8.98755*(10^9)*9.1/(11.3)^2= 1.84696*109

2) zero